create new tag
view all tags

20100603 Shift Log

First half hour getting beam into damping ring.

Laser Slave 1 failed again. Needed restarting.

Loaded our electron beam optics. Background around 500mW. Wire scanner (MW3X) touching beam again. Background now 120mV. Still something registering on the PLIC loss monitor.

BPM Data saved 10jun03_103346. 100mV background. No PLIC loss monitor signal. Beam relatively flat at our location. Using laser-wire optics.

Taking access to change to OTR. Rotated mirror after recollimating lens on laser-wire beam line in the hope that it would improve the transmission of both the OTR. This is because it is a standard 3" green mirror and is highly polarisation sensitive.

Used OTR to image beam.

Vertical 13 microns Horizontal 40 microns

Used hole burnt by laser to align chamber horizontally. Below in figures 1 and 2 the profile of the screen was taken by looking at the signal in the wire scanner detector in front of our detector. This was used instead of our own detector as the signal would saturate our PMT.


Figure 1: Wire scanner detector signal with screen around 50 microns into electron beam.


Figure 2: Wire scanner detector signal with screen around 250 microns into electron beam.

Figure 1 shows lots of notches / bumps in the edge of the screen, but Figure 2 shows one clear notch corresponding to the notch carved by the laser on a previous shift. This was centred on a chamber position of 1700. The chamber was moved here.

After this, we moved the screen vertically downwards to find out where it cut of the laser beam. Although the notch it created would seem to be around 500 microns wide, the top cut off was quite clear. This was found at a screen position of 9510. The screen was then moved upwards to unblock the laser by 20 - 50 microns (???).

The chamber was then moved down until the wire scanner detector signal was halved. This was found at a chamber position of 550 microns (vertically). Not by the dump charge monitor as usual.

After this, a timing scan was performed. However, the new oscilloscope produced very weird signals. It was observed that the BPM in one channel was showing up in all channels despite there being nothing plugged into them. After several reboots and calibrations, a situation was eventually found where the signal was partially visible, although it had a sinusoidal modulation on it. The original scope with a 1Ghz bandwidth screen is not working and so it cannot be used. We had to give the other scope back and it's no use. We now do not have a scope.

We got the timing within around 400ps with that scope at the end though.

Did a vertical chamber scan from 350 to 750 microns at a phase shifter voltage of 3.6V. Observed nothing.

Temporal scan from 1.5V to 5.5V for each vertical position starting at 545 going up (down in real space) 710 in 5 micron steps. Nothing observed

Finally, phase shifter set to 3.6V. Vertical scan at 2 microns per second from 715 to 1215. Thought we saw a bump at 925.

Scan from 900 to 950 at 1 micron per second, but nothing observed.

Had 30 minute extension to the shift for the last couple of scans.

Camac crashed when we tried to use the phase scanner so had to do it manually.

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r2 - 03 Jun 2010 - LaurieNevay

This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2024 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback